Arieonna:

2/10 MATH please help me

NotLeiland:

I gotchu

Arieonna:

Arieonna:

wait how A

Arieonna:

@laylalyssa @jhonyy9

NotLeiland:

the square root of 100 is 10...

NotLeiland:

It's A

thatgirlcrazy:

yea he know's what he's going cause 10 ^2 is 100 and the square root of 100 is 10 so a

Arieonna:

it was F

thatgirlcrazy:

how was it f

thatgirlcrazy:

oh wait I see how it was f I just don't know how to explain it

jhonyy9:

\[x^{2} = 100 => \sqrt{100} = +/- 10 \] because (-10)×(-10) = 100 and 10×10 = 100